use the information provided to write the standard form equation of each hyperbola
x^2-y^2+20x+19=0

Answer by stanbon(75887)

use the information provided to write the standard form equation of each hyperbola
x^2-y^2+20x+19=0
----
x^2+20x+? - y^2 = -19
-------
Complete the square::
x^2 + 20x+100 - y^2 = -19+100
-----
Factor:
(x+10)^2 - (y-0)^2 = 81
----------------
Cheers,
Stan H.
------------